OpenStudy (anonymous):

Solve the equation: . Round the solution to the nearest tenth. A. 1.7 B. 2.0 C. 3.2 D. 1.2

sam (.sam.):

Wheres the question

OpenStudy (anonymous):

oh sorry my bad. Give me a sec.

OpenStudy (anonymous):

x/3+x/4=1

OpenStudy (anonymous):

@.Sam.

sam (.sam.):

Common denominator

OpenStudy (anonymous):

12

sam (.sam.):

What do you get for numerator?

sam (.sam.):

\[\frac{x}{3}+\frac{x}{4}=1 \\ \\ \frac{4x+3x}{12}=1\]

OpenStudy (anonymous):

wait never mind. I think I got it. Is is 2.0?

sam (.sam.):

nope

sam (.sam.):

\[\frac{4x+3x}{12}=1 \] \[7x=12\] \[x=?\]

OpenStudy (anonymous):

1.7

sam (.sam.):

Yess

OpenStudy (anonymous):

Oh, I see. Thank you very much!

sam (.sam.):

yw :)